OpenStudy (anonymous):

Please help?? f(x) = √ (x+8) - x - 1 ; f(√ (2)+1) = ?

OpenStudy (anonymous):

\[f(\sqrt{3}) =(\sqrt{(3)}+8)-(\sqrt{3})-1 \] I think this is supposed to be your setup.

OpenStudy (anonymous):

\[f(\sqrt{2}+1) =((\sqrt{(2)}+1)+8)-(\sqrt{2}+1)-1\] or perhaps this

OpenStudy (anonymous):

I think I tried that. None of my answers are coming out right though.